# LieRE: Lie Rotational Positional Encodings

### v3

@@ Appendix B Experiment Hyperparameters
B.3 Multi-resolution Classification
The second training recipe consists of 30 epochs with an initial learning rate of 1E-5 with a cosine decay. This mirrors the DEIT III training reciple that first pretrains at a lower resolution and finetunes at a higher resolution.
B.4 CIFAR-100 Performance Across Model Scales
We also evaluate how the inclusion of LieRE affects performance across model sizes on the CIFAR-100 in table 5. We observe that LieRE retains a statistically significant lead in performance across all three model sizes.
B.4.1 Basis parameters
@@
B.5 Python implementation of LieRE rotation matrix computation
1basis_raw_params = nn.Parameter(
2 torch.rand(
3 input_dimensionality,
4 head_dim,
5 head_dim,
6 ) * 2 * math.pi # optional, inspired from RoPE-Mixed paper
7)
8
9upper_triangle = (
10 torch.triu(basis_raw_params, diagonal=1)
11)
12skew_bases = upper_triangle - torch.transpose(upper_triangle, -1, -2)
13in_basis_positions = (
14 positions.reshape(list(positions.shape) + [1] * 2) * skew_bases
15)
16rotation_log = torch.sum(in_basis_positions, dim=-3)
17rotation = torch.matrix_exp(rotation_log.to(dtype=torch.float32)).to(dtype=positions.dtype)
B.6 Compute efficiency
@@
B.7 FLOPS Comparison of methods
@@
B.8 Validation Losses
Table 7: 2D image and 3D video classification Top-1 Validation loss (95% confidence intervals) results. All models use 85.1M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Stein et al., 2019) equivalent to DeiT, ∗∗ equivalent to Vivit (spatio-temporal).
@@ B.8 Validation Losses
B.9 Basis parameters
@@
B.10 Validation Accuracy Curves
Figure 11: CIFAR-100
B.11 Training Loss Curves
Figure 12: CIFAR-100
B.12 Attention Maps

### v4

@@ B.1 Experimental Hyperparameters
B.4 Multi-resolution Classification
The second training recipe consists of 30 epochs with an initial learning rate of 1E-5 with a cosine decay. This mirrors the DEIT III training recipe that first pretrains at a lower resolution and finetunes at a higher resolution.
B.5 CIFAR-100 Performance Across Model Scales
We further evaluate the impact of incorporating LieRE across different model sizes on CIFAR-100, as shown in Table 6. LieRE consistently outperforms the baseline with statistically significant gains across all three model scales. However, these results may be sensitive to dataset size, as all models are trained from scratch in this study. This is reflected in the performance drop observed with the ViT-Huge model.
@@ B.5 CIFAR-100 Performance Across Model Scales
B.6 Initialization Sensitivity
In order to understand the sensitivity of the methods to the initialization, we explore several scaling factor for the weights initialization. The results are presented in table 7.
Table 7: Paired Z-test between $2\pi$ and $1$ initialization for LieRE8 and RoPE-Mixed.
Metric LieRE8 ($2\pi$ vs 1 init) RoPE-Mixed ($2\pi$ vs 1 init)
Z-statistic -1.81 -3.85
P-value 0.070 0.00012
Difference between means -0.0118 -0.0255
95% Confidence Interval [-0.0246, 0.0010] [-0.0385, -0.0125]
Appendix C Python implementation of LieRE rotation matrix computation
basis_raw_params = nn.Parameter( torch.rand( input_dimensionality, head_dim, head_dim, ) * 2 * math.pi # optional, inspired from RoPE-Mixed paper ) upper_triangle = ( torch.triu(basis_raw_params, diagonal=1) ) skew_bases = upper_triangle - torch.transpose(upper_triangle, -1, -2) in_basis_positions = ( positions.reshape(list(positions.shape) + [1] * 2) * skew_bases ) rotation_log = torch.sum(in_basis_positions, dim=-3) rotation = torch.matrix_exp(rotation_log.to(dtype=torch.float32)) rotation = rotation.to(dtype=positions.dtype)
Appendix D Compute efficiency
@@
D.1 FLOPS Comparison of methods
@@
Appendix E Validation Losses
Table 9: 2D image and 3D video classification Top-1 Validation loss (95% confidence intervals) results. All models use 85.1M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Flanders et al., 2020) equivalent to DeiT, ∗∗ equivalent to Vivit (spatio-temporal).
@@ Appendix E Validation Losses
Appendix F Basis parameters scaling
@@
Appendix G Attention Maps